OpenStudy (anonymous):

Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. (3 points) Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. (3 points) Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work? (4 points)

OpenStudy (texaschic101):

A.) W = 10h -- h stands for hrs worked B.) not sure C.) 40 hrs regular pay = 40 * 10 = 400 that leaves $90 overtime pay 90/15 = 6 so he worked 10 hrs regular pay + 6 hrs overtime pay